# Deep Multi-Task Learning Has Low Amortized Intrinsic Dimensionality

### v1

@@ Appendix B Experimental Details
B.1 Model Training Details
As mentioned in the section 7, we use an MLP model for Products and Folktables datasets, and convolutional networks for MNIST and CIFAR datasets. The details of the model architectures are provided in Table 5.
@@ B.1 Model Training Details
The detailed numeric values used for quantizing parameters in Table 2 are shown in Table 4.
@@
Single Task ^$\widehat{\mathcal{R}}(f)$ [%] 23.4 19.4 28.0 16.0 30.4 63.9
codebook size $r$ 10 10 5 5 10 10
(average) encoding length 855.4 854.3 212.6 216.4 542.5 542.0
Multi-task ^$\widehat{\mathcal{R}}(f_{1},\dots,f_{n})$ [%] 10.1 6.6 27.2 13.9 30.5 62.7
codebook sizes $r_{g}$ / $r_{l}$ 10 / 3 15 / 10 10 / 10 10 / 10 10 / 20 20 / 10
$l(E)$ 2323 14887 1586 1192 2358 4209
$l_{E}(f_{1},\dots,f_{n})$ 508 4796 686 1128 4144 3341
(average) encoding length 94.4 651.1 37.9 38.7 65.0 75.5
@@
split-CIFAR10/ split-CIFAR100 Conv1 Conv2d(input: $C$, output: $16$, kernel: $5\times 5$)
@@
Products/ Folktables FC1 Linear(input: input_dim, output: $128$)
Activation ReLU
FC2 Linear(input: $128$, output: $64$)
Activation ReLU
FC3 Linear(input: $64$, output: $32$)
Output Linear(input: $32$, output: Output_dim)

### v2

@@ Appendix B Experimental Details
B.1 Datasets
MNIST Shuffled Pixels (SP): (Amit & Meir, 2018) each task is a random subset of the MNIST (LeCun & Cortes, 1998) dataset in which 200 of the input pixels are randomly shuffled. The same shuffling is consistent across all samples of that task.
Folktables: (Ding et al., 2021a) A tabular dataset consisting of public US census information. From personal features, represented in a binary encoding, the model should predict if a person’s income is above or below a threshold. Tasks correspond to different geographic regions.
B.2 Model architectures
For the MNIST experiments, we use convolutional networks used in Amit & Meir (2018). For the vectorial dataset Product and the tabular dataset Folktables, we use 4-layer fully connected networks. For the CIFAR experiments, we use two different networks: 1) The CNN used in Scott et al. (2024), and a ViT model pretrained from ImageNet (Dosovitskiy et al., 2021). The details of the model architectures are provided in Table 6.
The models’ ambient dimensions (number of network weights) range from approximately 12000 to approximately 5.5 million, i.e. far more than the available number of samples per task. As random matrices $P$ for the single-task parametrization (1), we use the Kronecker product projector of Lotfi et al. (2022), $P=Q_{1}\otimes Q_{2}/\sqrt{D}$, for $Q_{1},Q_{2}\sim\mathcal{N}(0,1)^{\sqrt{D}\times\sqrt{d}}$. By this construction, the matrix $P\in\mathbb{R}^{D\times d}$ never has to be explicitly instantiated, which makes the memory and computational overhead tractable. For the multi-task representation (3), we use the analogous construction to form $Q^{\prime}=[P_{1},P_{2},\dots,P_{k}]=Q^{\prime}_{1}\otimes Q^{\prime}_{2}/\sqrt{D}\in\mathbb{R}^{D\times kl}$.
@@ B.2 Model architectures
B.3 Model training details
@@
The detailed numeric values used for quantizing parameters in Table 2 are shown in Table 4 and Table 5.
Table 4: Numeric values contributing to the generalization bounds in Table 2 for $n$: number of tasks, $m$: average number of examples per task, $L$: number of classes.
dataset MNIST SP MNIST PL Folktables Products
$n$ / $m$ / $L$ 30 / 2000 / 10 30 / 2000 / 10 60 / 900 / 2 60 / 2000 / 2
Single Task ^$\widehat{\mathcal{R}}(f)$ 0.234 0.194 0.280 0.160
codebook size $r$ 10 10 5 5
(average) encoding length 855.4 854.3 212.6 216.4
Multi-task ^$\widehat{\mathcal{R}}(f_{1},\dots,f_{n})$ 0.101 0.066 0.272 0.139
codebook sizes $r_{g}$ / $r_{l}$ 10 / 3 15 / 10 10 / 10 10 / 10
$l(E)$ 2323 14887 1586 1192
$l_{E}(f_{1},\dots,f_{n})$ 508 4796 686 1128
(average) encoding length 94.4 651.1 37.9 38.7
@@
dataset split-CIFAR10 split-CIFAR100
Model CNN ViT CNN ViT
$n$ / $m$ / $L$ 100 / 453 / 3 30 / 1248 / 3 100 / 450 / 10 30 / 1250 / 10
Single Task ^$\widehat{\mathcal{R}}(f)$ 0.304 0.181 0.639 0.314
codebook size $r$ 10 10 10 10
(average) encoding length 542.5 861.2 542.0 1842.9
Multi-task ^$\widehat{\mathcal{R}}(f_{1},\dots,f_{n})$ [%] 0.305 0.106 0.627 0.274
codebook sizes $r_{g}$ / $r_{l}$ 10 / 20 10 / 20 20 / 10 20 / 30
$l(E)$ 2358 3109 4209 11512
$l_{E}(f_{1},\dots,f_{n})$ 4144 1747 3341 4957
(average) encoding length 65.0 161.9 75.5 549.0
@@
Products/ Folktables FC1 Linear(input: input_dim, output: $128$)
Activation ReLU
FC2 Linear(input: $128$, output: $64$)
Activation ReLU
FC3 Linear(input: $64$, output: $32$)
Output Linear(input: $32$, output: Output_dim)
split-CIFAR10/ split-CIFAR100 (ConvNet) Conv1 Conv2d(input: $C$, output: $16$, kernel: $5\times 5$)
@@
split-CIFAR10/ split-CIFAR100 (ViT) Patch Embedding Conv2d(input: $3$, output: $192$, kernel: $16\times 16$, stride: $16$)
12 Transformer Blocks
LayerNorm(shape: $192$, eps: $1\text{e}{-6}$)
Attention:
   Linear(input: $192$, output: $576$) for qkv
   Linear(input: $192$, output: $192$) for projection
MLP:
   LayerNorm(shape: $192$, eps: $1\text{e}{-6}$)
   Linear(input: $192$, output: $768$)
   Activation: GELU
   Linear(input: $768$, output: $192$)
Post-Norm LayerNorm(shape: $192$, eps: $1\text{e}{-6}$)
Classification Head Linear(input: $192$, output: Output_dim)